OpenStudy (annie96):

Allan and Dave bowl together and their combined total score for one game was 375 points. Allan’s score was 60 less than twice Dave’s. What were their scores? Which is a system of equations to model the problem if x represents Dave’s score and y represents Allan’s score? (Points : 1) x + y = 60 y = 2x – 375 x + y = 375 y = 2x – 60 x + y = 375 y = 2x + 60 x – y = 375 y = 2x – 60

OpenStudy (mathstudent55):

Since their added scores are 375, and one is x and the other is y, then you know x + y = 375. You can eliminate the 4th choice.

OpenStudy (mathstudent55):

We need to write this as an equation: "Allan’s score was 60 less than twice Dave’s. "

OpenStudy (annie96):

So D.

OpenStudy (mathstudent55):

Did you read what I just wrote? We can eliminate D since instead of having an equation x + y = 375, it has x - y = 375. So far all we know is that D is definitely not it.

OpenStudy (mathstudent55):

Now we need to write "Allan’s score was 60 less than twice Dave’s. " as an equation. We are told Allan's score is y and Dave's score is x. Let's do that below the statement. "Allan’s score was 60 less than twice Dave’s. " y's score was 60 less than twice x's y = 2x - 60

OpenStudy (mathstudent55):

The second equation is y = 2x - 60

OpenStudy (annie96):

x + y = 375 y = 2x – 60

OpenStudy (mathstudent55):

correct